A normally distributed population has a mean of 25.6 and a standard deviation of 3.3. Find the probability that a single randomly selected element X of the population exceeds 30.



1
Expert's answer
2022-06-02T06:25:38-0400

We have a normal distribution, "\mu=25.6, \sigma=3.3."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}=\cfrac{30-25.6}{3.3}=1.33;"

"P(X>30)=P(Z>1.33)=1-P(Z<1.33)=\\\\\n=1-0.9082=0.0918\\ \\text{\u200b(from z-table).}"